A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 2.80 s for the ball to reach its maximum height.
(a) Find its initial velocity.
m/s (upward)

(b) Find the height it reaches.
m
Answer #1
(a)
At the maximum height, the velocity is .


(b)
Answer #2

(a) By v = u -gt
=>0 = u - 9.8 x 2.80
=>u = 27.44 m/s
(b) By v^2 = u^2 - 2gh
=>0 = (27.44)^2 - 2 x 9.8 x h
=>h = 38.42 m